the manager of a football team requires 500 soccer balls for training purposes and donatiuons to disadvantaged communities. he comes across four suppliers fifa, uefa, cosafa and jaffer that sells these soccer balls. fifa sells balls at R140.00 each, while uefa sells the balls aT r180.00 each with a 20% discount on the total cost. Cosafa have a buy one get one free special on the balls, with the cost of each ball being R300.00, while JAFFER sells each ball at R220.00 with the third off the total cost. Which supplier offers the best deal? Show all calculations in detail.
Answer by jorel555(1290)

Let's figure out the total cost of each:
fifa: 500 x 140=70000
uefa: 500 x 180=90000 - 20%= 72000
cosafa 500 x 300=150000 div. by 2=75000
JAFFER: 500 x 220=110000 - 1/3=73333.333
Your best deal is with fifa!!!!!!!!!